what dimensions should the rancher use to obtain the area of 350 with 100 meters of fence?"
Answer by ankor@dixie-net.com(22740)

what dimensions should the rancher use to obtain the area of 350 with 100 meters of fence?"
:
Using the perimeter equation, get L in terms of W
2L + 2W = 100
Simplify, divide by 2
L + W = 50
L = (50-W)
:
The area equation
L * W = 350
Substitute (50-W) for L
(50-W) * W = 350
-W^2 + 50 W = 350
-W^2 + 50W - 350 = 0
W^2 - 50W + 350 = 0
Will not factor use the quadratic formula:
a=1, b=-50; c=350
:
You should get the lesser solution: W = 8.416 is the width
then (the greater solution)
L = 50 - 8.416 = 41.584 is the length